# Enhancing Fairness in Unsupervised Graph Anomaly Detection through Disentanglement

### v1

@@ 3.2. Disentangled Fair Representation Learning
The encoder $f_{e}$ is designed to learn disentangled node representations, premised on the assumption that the latent space can be decomposed into two independent subspaces: one associated with sensitive attributes and the other independent of sensitive attributes. For the graph data, the posterior distribution of node representation $q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})$ is based on the node attributes $\mathbf{X}$ and graph topology $\mathbf{A}$. To ensure disentanglement on graphs, it is essential to establish conditional independence between sensitive-irrelevant representations $\mathbf{Z}_{x}$ and sensitive-relevant representations $\mathbf{Z}_{s}$, given $\mathbf{X}$ and $\mathbf{A}$. Under the assumption of conditional independence, the variational posterior distribution $q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})$ can be defined as the product of the individual distributions for $\mathbf{Z}_{x}$ and $\mathbf{Z}_{s}$:
(3) $q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})=q(\mathbf{Z}_{x}|\mathbf{X},\mathbf{A})q(\mathbf{Z}_{s}|\mathbf{X},\mathbf{A}).$
@@
As the disentanglement is conducted in the latent space, it is crucial to capture informative node representations from both node attributes $\mathbf{X}$ and topological structure $\mathbf{A}$. Thus, we adopt a GNN as the backbone of the encoder $f_{e}$. In this work, we take Graph Convolutional Network (GCN) (Kipf and Welling, 2016) as an example. For the $l$-th convolutional layer, the node representation $\mathbf{H}^{(l)}$ is updated by:
(4) $\mathbf{H}^{(l)}=\text{Conv}(\mathbf{H}^{(l-1)},\mathbf{A})=\phi\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l-1)}\mathbf{W}^{l}\right),$
@@
(5) $\mathbf{Z}_{x}=\bm{\mu}+\bm{\sigma}\cdot\bm{\varepsilon},\ \bm{\varepsilon}\sim\mathcal{N}(\mathcal{\mathbf{0},\mathbf{I}}),$
where $\mathcal{N}(\mathcal{\mathbf{0},\mathbf{I}})$ is the standard Gaussian distribution with a mean vector of zeros $\mathbf{0}$ and an identity covariance matrix $\mathbf{I}$. Here, the mean matrix $\bm{\mu}=\text{Conv}_{\bm{\mu}}(\text{Conv}_{\text{shr}}(\mathbf{X},\mathbf{A}),\mathbf{A})$ and the log standard deviation matrix $\log\bm{\sigma}=\text{Conv}_{\bm{\sigma}}(\text{Conv}_{\text{shr}}(\mathbf{X},\mathbf{A}),\mathbf{A})$ are obtained by GCNs with a shared convolutional layer $\text{Conv}_{\text{shr}}$. On the other hand, we estimate the sensitive-relevant representations $\mathbf{Z}_{s}$ in a deterministic manner:
(6) $\mathbf{Z}_{s}=\text{Conv}_{\bm{\varphi}}(\text{Conv}_{\text{shr}}(\mathbf{X},\mathbf{A}),\mathbf{A}),$
@@
In the decoding phase of disentangled representation learning, we reconstruct the attribute matrix $\mathbf{X}$, the adjacency matrix $\mathbf{A}$, and sensitive attributes $\mathbf{S}$ from the latent space. Due to sensitive-irrelevant representations $\mathbf{Z}_{x}$ are conditionally independent of the given sensitive attributes $\mathbf{S}$, the probability $p(\mathbf{S}|\mathbf{Z}_{x},\mathbf{Z}_{s})$ can be simplified to $p(\mathbf{S}|\mathbf{Z}_{s})$. Thus, the reconstruction phase in DEFEND can be formulated by:
(7) $p(\mathbf{X},\mathbf{A},\mathbf{S}|\mathbf{Z}_{x},\mathbf{Z}_{s})=p(\mathbf{X}|\mathbf{Z}_{x},\mathbf{Z}_{s})p(\mathbf{A}|\mathbf{Z}_{x},\mathbf{Z}_{s})p(\mathbf{S}|\mathbf{Z}_{s}).$
@@
To model the distribution of reconstructed node attributes $p(\mathbf{X}|\mathbf{Z}_{x},\mathbf{Z}_{s})$, we adopt an attribute decoder $f_{x}$, which has the reverse structure of the encoder $f_{e}$. Furthermore, we consider an inner product decoder $f_{a}$ for structure reconstruction as follows:
(8) $p(\mathbf{A}|\mathbf{Z}_{x},\mathbf{Z}_{s})=\text{Sigmoid}(\mathbf{Z}\mathbf{Z}^{\top}),$
@@
where $\mathbf{Z}=\text{Concat}(\mathbf{Z}_{x},\mathbf{Z}_{s})$. Next, to model the conditional probability distribution $p(\mathbf{S}|\mathbf{Z}_{s})$, we adopt a binary classifier, where sensitive attribute $\mathbf{S}$ is assumed to follow a Bernoulli distribution parameterized by the Sigmoid function applied to $\mathbf{Z}_{s}$. The conditional distribution for the sensitive attributes is formulated as:
(9) $p(\mathbf{S}|\mathbf{Z}_{s})=\text{Bernoulli}(\mathbf{S}|\text{Sigmoid}(\mathbf{Z}_{s})).$
3.2.3. Adversary.
@@
To encourage the independence between sensitive-irrelevant representations $\mathbf{Z}_{x}$ and sensitive-relevant representations $\mathbf{Z}_{s}$, it is imperative that the aggregate posterior distribution can be factorized as $q(\mathbf{Z}_{x},\mathbf{Z}_{s})=q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})$. Thus, we employ the Kullback-Leibler (KL) divergence between $q(\mathbf{Z}_{x},\mathbf{Z}_{s})$ and $q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})$ as the disentanglement criteria, which can be expressed as ${KL}\big{[}q(\mathbf{Z}_{x},\mathbf{Z}_{s})\,\|\,q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})\big{]}$. Following previous studies (Kim and Mnih, 2018; Sugiyama et al., 2012; Creager et al., 2019), we adopt a binary adversary to encourage the disentanglement by approximating the log density ratio inherent in the KL divergence term:
(10) ${KL}\big{[}q(\mathbf{Z}_{x},\mathbf{Z}_{s})\,\|\,q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})\big{]}=\mathbb{E}_{q(\mathbf{Z}_{x},\mathbf{Z}_{s})}\log\frac{q(\mathbf{Z}_{x},\mathbf{Z}_{s})}{q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})}$
$\approx\mathbb{E}_{q(\mathbf{Z}_{x},\mathbf{Z}_{s})}[\log g_{\omega}(\tilde{\mathbf{y}}=1|\mathbf{Z}_{x},\mathbf{Z}_{s})-\log g_{\omega}(\tilde{\mathbf{y}}=0|\mathbf{Z}_{x},\mathbf{Z}_{s})],$
@@
where $\tilde{\mathbf{y}}=1$ denotes ”true” samples from the aggregate posterior $q(\mathbf{Z}_{x},\mathbf{Z}_{s})$ and $\tilde{\mathbf{y}}=0$ denotes ”fake” samples from the product of the marginals $q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})$. Specifically, we implement an MLP as the binary adversary $g_{\omega}$ to predict whether $\mathbf{z}_{x}^{i}$, the sensitive-irrelevant representation of node $v_{i}$, and $\mathbf{z}_{s}^{j}$, the sensitive-relevant of node $v_{j}$, are from the same node.
3.2.4. Learning Objective.
@@
In the disentangled representation learning phase, we optimize the disentangled graph encoder $f_{e}$, structure decoder $f_{a}$, and attribute decoder $f_{x}$ from three aspects, including the variational lower bound term, the disentanglement term, and the predictiveness term. First, the variational lower bound term comprises a reconstruction term and a KL divergence regularization term. The reconstruction term, addressing both node attributes and topological structure, can be formulated as:
(11) $\mathcal{L}_{rec}=\mathbb{E}_{q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})}\left[(1-\epsilon)\log p(\mathbf{X}|\mathbf{Z}_{x},\mathbf{Z}_{s})+\epsilon\log p(\mathbf{A}|\mathbf{Z}_{x},\mathbf{Z}_{s})\right]$
@@
where the first term is the reconstruction error for node attribute and the second term is the reconstruction error for graph topology. Moreover, $\epsilon=\frac{\sigma_{\mathbf{X}}}{\sigma_{\mathbf{X}}+\sigma_{\mathbf{A}}}$ is the weight coefficient for balancing the impact of structure and attribute reconstruction, where $\sigma_{\mathbf{A}}$ and $\sigma_{\mathbf{X}}$ denotes the standard deviations of $\mathbf{A}$ and $\mathbf{X}$, respectively. The KL divergence regularization term is employed to minimize the KL divergence between the posterior distribution $q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})$ and the prior distribution $p(\mathbf{Z}_{x},\mathbf{Z}_{s})$. The overall objective within the variational lower bound can be formulated as:
(12) $\mathcal{L}_{vae}=\mathcal{L}_{rec}-KL[q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})\parallel p(\mathbf{Z}_{x},\mathbf{Z}_{s})],$
@@
where $p(\mathbf{Z}_{x},\mathbf{Z}_{s})=p(\mathbf{Z}_{x})p(\mathbf{Z}_{s})$ under the assumption that $\mathbf{Z}_{x}$ and $\mathbf{Z}_{s}$ are independent. The prior distributions $p(\mathbf{Z}_{x})$ and $p(\mathbf{Z}_{s})$ are modeled by the standard Gaussian distribution and uniform distribution, respectively. Next, the disentanglement term, which is crucial for separating sensitive-relevant and sensitive-irrelevant representations, is calculated by the binary adversary $g_{\omega}$:
(13) $\mathcal{L}_{dis}=\mathbb{E}_{\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i}\sim q(\mathbf{Z}_{x},\mathbf{Z}_{s})}\log p(\tilde{y}=1|\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i})-\log p(\tilde{y}=0|\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i}),$
@@
where $(\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i})$ is a true sample for which $\tilde{y}=1$. Moreover, DEFEND incorporates a predictiveness term to align the sensitive representation $\mathbf{Z}_{s}$ closely with the given sensitive attributes $\mathbf{S}$. Intuitively, accurate prediction of sensitive attributes can enhance the comprehensive understanding of node attributes in latent space and facilitate a clearer distinction between representations associated with sensitive attributes and those that are not. The predictiveness loss term can be expressed as:
(14) $\mathcal{L}_{pre}=\mathbb{E}_{q(\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})}\log p(\mathbf{S}|\mathbf{Z}_{s}).$
@@
(15) $\mathcal{L}_{total}=\mathcal{L}_{vae}+\gamma\mathcal{L}_{dis}+\alpha\mathcal{L}_{pre},$
@@
During the training phase of the binary adversary $g_{\omega}$, the true sample $(\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i})$ is sampled from the aggregate posterior $q(\mathbf{Z}_{x},\mathbf{Z}_{s})$ while the fake sample $(\mathbf{z}_{x}^{j},\mathbf{z}_{s}^{k})$ is sampled from the product of marginal posterior distributions $q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})$. The adversarial loss can be formulated as:
(16) $\mathcal{L}_{adv}$ $=\mathbb{E}_{\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i}\sim q(\mathbf{Z}_{x},\mathbf{Z}_{s})}\log p(\tilde{y}=1|\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i})$
$+\mathbb{E}_{\mathbf{z}_{x}^{j},\mathbf{z}_{s}^{k}\sim q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})}\log[1-p(\tilde{y}=0|\mathbf{z}_{x}^{j},\mathbf{z}_{s}^{k})].$
@@
In DEFEND, the optimization of $f_{e}$, $f_{a}$ and $f_{x}$ and that of $g_{\omega}$ are conducted adversarially. A well-trained encoder $f_{e}$, which can effectively encode node attributes and its local topology into latent space, and obtain sensitive-irrelevant representations, can be adeptly employed in downstream tasks for enhancing fairness.
3.3. Graph Anomaly Detection

### v2

@@ 3.1.1 Disentangled Graph Encoder
Based on the assumption that the latent space can be decomposed into two independent subspaces (Creager et al., 2019; Oh et al., 2022): one associated with sensitive attributes and the other devoid of them, the disentangled graph encoder $f_{e}$ strives to capture informative node representations that are irrelevant to sensitive attributes. For graph data, the posterior distribution of node representations $q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})$ is derived from node attributes $\mathbf{X}$ and graph topology $\mathbf{A}$. To achieve disentanglement, it is necessary to establish conditional independence between sensitive-irrelevant representations $\mathbf{Z}_{x}$ and sensitive-relevant representations $\mathbf{Z}_{s}$, given $\mathbf{X}$ and $\mathbf{A}$. The variational posterior distribution $q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})$ can be defined as the product of the individual distributions for $\mathbf{Z}_{x}$ and $\mathbf{Z}_{s}$:
$q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})=q(\mathbf{Z}_{x}|\mathbf{X},\mathbf{A})q(\mathbf{Z}_{s}|\mathbf{X},\mathbf{A}).$ (3)
@@
As the disentanglement is conducted in latent space, it is crucial to capture informative node representations from both node attributes and topological structure. Thus, we adopt GNNs as the backbone of the disentangled graph encoder $f_{e}$. In this work, we take Graph Convolutional Network (GCN) (Kipf & Welling, 2016) as an example. For the $l$-th convolutional layer, the node representation $\mathbf{H}^{(l)}$ is updated by:
$\mathbf{H}^{(l)}=\text{Conv}(\mathbf{H}^{(l-1)},\mathbf{A})=\phi\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l-1)}\mathbf{W}^{l}\right),$ (4)
@@
$\mathbf{Z}_{x}=\bm{\mu}+\bm{\sigma}\cdot\bm{\varepsilon},\ \bm{\varepsilon}\sim\mathcal{N}(\mathcal{\mathbf{0},\mathbf{I}}),$ (5)
where $\mathcal{N}(\mathcal{\mathbf{0},\mathbf{I}})$ is the standard Gaussian distribution with a mean vector of zeros $\mathbf{0}$ and an identity covariance matrix $\mathbf{I}$. The mean matrix $\bm{\mu}=\text{Conv}_{\bm{\mu}}(\text{Conv}_{\text{shr}}(\mathbf{X},\mathbf{A}),\mathbf{A})$ and the log standard deviation matrix $\log\bm{\sigma}=\text{Conv}_{\bm{\sigma}}(\text{Conv}_{\text{shr}}(\mathbf{X},\mathbf{A}),\mathbf{A})$ are obtained by GCNs with a shared convolutional layer $\text{Conv}_{\text{shr}}$. Following (Creager et al., 2019), we estimate the sensitive-relevant representations $\mathbf{Z}_{s}$ in a deterministic manner:
$\mathbf{Z}_{s}=\text{Conv}_{\bm{\varphi}}(\text{Conv}_{\text{shr}}(\mathbf{X},\mathbf{A}),\mathbf{A}).$ (6)
@@ 3.1.2 Decoders
In the decoding phase of disentangled representation learning, we reconstruct the attribute matrix $\mathbf{X}$, the adjacency matrix $\mathbf{A}$, and sensitive attributes $\mathbf{S}$. Due to sensitive-irrelevant representations $\mathbf{Z}_{x}$ are conditionally independent of the given sensitive attributes $\mathbf{S}$, the probability $p(\mathbf{S}|\mathbf{Z}_{x},\mathbf{Z}_{s})$ can be simplified to $p(\mathbf{S}|\mathbf{Z}_{s})$. Consequently, the decoding phase can be formulated by:
$p(\mathbf{X},\mathbf{A},\mathbf{S}|\mathbf{Z}_{x},\mathbf{Z}_{s})=p(\mathbf{X}|\mathbf{Z}_{x},\mathbf{Z}_{s})p(\mathbf{A}|\mathbf{Z}_{x},\mathbf{Z}_{s})p(\mathbf{S}|\mathbf{Z}_{s}).$ (7)
@@
We adopt an attribute decoder $f_{x}$ based on GNNs to model the distribution of reconstructed node attributes $p(\mathbf{X}|\mathbf{Z}_{x},\mathbf{Z}_{s})$. Besides, we consider an inner product decoder $f_{a}$ for structure reconstruction:
$p(\mathbf{A}|\mathbf{Z}_{x},\mathbf{Z}_{s})=\text{Sigmoid}(\mathbf{Z}\mathbf{Z}^{\top}),$ (8)
@@
where $\mathbf{Z}=\text{Concat}(\mathbf{Z}_{x},\mathbf{Z}_{s})$. Next, to model the conditional probability distribution $p(\mathbf{S}|\mathbf{Z}_{s})$, we employ a binary classifier, where sensitive attributes $\mathbf{S}$ are assumed to follow a Bernoulli distribution parameterized by the Sigmoid function applied to $\mathbf{Z}_{s}$. The conditional distribution for sensitive attributes is formulated as:
$p(\mathbf{S}|\mathbf{Z}_{s})=\text{Bernoulli}(\mathbf{S}|\text{Sigmoid}(\mathbf{Z}_{s})).$ (9)
3.1.3 Adversary
@@
To encourage the independence between sensitive-irrelevant representations $\mathbf{Z}_{x}$ and sensitive-relevant representations $\mathbf{Z}_{s}$, it is imperative that the aggregate posterior distribution can be factorized as $q(\mathbf{Z}_{x},\mathbf{Z}_{s})=q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})$. Thus, we employ the Kullback-Leibler (KL) divergence between $q(\mathbf{Z}_{x},\mathbf{Z}_{s})$ and $q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})$ as the disentanglement criteria, which can be expressed as ${KL}\big{[}q(\mathbf{Z}_{x},\mathbf{Z}_{s})\,\|\,q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})\big{]}$. Following (Kim & Mnih, 2018; Sugiyama et al., 2012; Creager et al., 2019), we adopt a binary adversary $g_{\omega}$ to encourage the disentanglement by approximating the log density ratio inherent in the KL divergence term:
${KL}\big{[}q(\mathbf{Z}_{x},\mathbf{Z}_{s})\,\|\,q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})\big{]}=\mathbb{E}_{q(\mathbf{Z}_{x},\mathbf{Z}_{s})}\log\frac{q(\mathbf{Z}_{x},\mathbf{Z}_{s})}{q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})}$ (10)
$\approx\mathbb{E}_{q(\mathbf{Z}_{x},\mathbf{Z}_{s})}[\log p(\tilde{\mathbf{y}}=1|\mathbf{Z}_{x},\mathbf{Z}_{s})-\log p(\tilde{\mathbf{y}}=0|\mathbf{Z}_{x},\mathbf{Z}_{s})],$
@@
where $\tilde{\mathbf{y}}=1$ denotes "true" samples from the aggregate posterior $q(\mathbf{Z}_{x},\mathbf{Z}_{s})$ and $\tilde{\mathbf{y}}=0$ denotes "fake" samples from the product of the marginals $q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})$. Specifically, we implement an MLP as $g_{\omega}$ to predict whether the sensitive-irrelevant representation $\mathbf{z}_{x}^{i}$ and the sensitive-relevant representation $\mathbf{z}_{s}^{j}$ are from the same node.
3.1.4 Learning Objective
@@
We optimize the disentangled graph encoder $f_{e}$, structure decoder $f_{a}$, and attribute decoder $f_{x}$ from three aspects, including the variational lower bound term, disentanglement term, and predictiveness term. First, the variational lower bound term comprises a reconstruction term and a KL divergence regularization term. The reconstruction term is defined as:
$\mathcal{L}_{rec}=\mathbb{E}_{q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})}\left[(1-\epsilon)\log p(\mathbf{X}|\mathbf{Z}_{x},\mathbf{Z}_{s})+\epsilon\log p(\mathbf{A}|\mathbf{Z}_{x},\mathbf{Z}_{s})\right],$ (11)
@@
where the first term is the reconstruction error for node attributes and the second term is for graph topology. Moreover, $\epsilon=\frac{\sigma_{\mathbf{X}}}{\sigma_{\mathbf{X}}+\sigma_{\mathbf{A}}}$ is the weight coefficient for automated balancing the impact of structure and attribute reconstruction (Liu et al., 2022b), where $\sigma_{\mathbf{A}}$ and $\sigma_{\mathbf{X}}$ denotes the standard deviations of $\mathbf{A}$ and $\mathbf{X}$, respectively. The KL divergence regularization term (Kingma, 2013) is employed to minimize the KL divergence between the posterior distribution $q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})$ and the prior distribution $p(\mathbf{Z}_{x},\mathbf{Z}_{s})$. Thus, the variational lower bound term can be formulated as:
$\mathcal{L}_{vae}=\mathcal{L}_{rec}-KL[q(\mathbf{Z}_{x},\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})\parallel p(\mathbf{Z}_{x},\mathbf{Z}_{s})],$ (12)
@@
where $p(\mathbf{Z}_{x},\mathbf{Z}_{s})=p(\mathbf{Z}_{x})p(\mathbf{Z}_{s})$ under the assumption that $\mathbf{Z}_{x}$ and $\mathbf{Z}_{s}$ are independent. The prior distributions $p(\mathbf{Z}_{x})$ and $p(\mathbf{Z}_{s})$ are modeled by the standard Gaussian distribution and uniform distribution, respectively. Next, the disentanglement term, which encourages the separation of sensitive-relevant and -irrelevant representations as detailed in Eq. (10), is calculated as follows:
$\mathcal{L}_{dis}=\mathbb{E}_{\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i}\sim q(\mathbf{Z}_{x},\mathbf{Z}_{s})}\log p(\tilde{y}=1|\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i})-\log p(\tilde{y}=0|\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i}).$ (13)
@@
Intuitively, accurate prediction of sensitive attributes can enhance the comprehensive understanding of node attributes in latent space and facilitate a clearer distinction between representations associated with sensitive attributes and those that are not. Thus, DEFEND incorporates a predictiveness term to align the sensitive representation $\mathbf{Z}_{s}$ closely with the given sensitive attributes $\mathbf{S}$:
$\mathcal{L}_{pre}=\mathbb{E}_{q(\mathbf{Z}_{s}|\mathbf{X},\mathbf{A})}\log p(\mathbf{S}|\mathbf{Z}_{s}).$ (14)
@@
$\mathcal{L}_{total}=\mathcal{L}_{vae}+\gamma\mathcal{L}_{dis}+\alpha\mathcal{L}_{pre},$ (15)
@@
To train the binary adversary $g_{\omega}$, the true sample $(\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i})$ is sampled from the aggregate posterior $q(\mathbf{Z}_{x},\mathbf{Z}_{s})$ while the fake sample $(\mathbf{z}_{x}^{j},\mathbf{z}_{s}^{k})$ is sampled from the product of marginal posterior distributions $q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})$. The adversarial loss is formulated as:
$\mathcal{L}_{adv}=\mathbb{E}_{\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i}\sim q(\mathbf{Z}_{x},\mathbf{Z}_{s})}\log p(\tilde{y}=1|\mathbf{z}_{x}^{i},\mathbf{z}_{s}^{i})+\mathbb{E}_{\mathbf{z}_{x}^{j},\mathbf{z}_{s}^{k}\sim q(\mathbf{Z}_{x})q(\mathbf{Z}_{s})}\log[1-p(\tilde{y}=0|\mathbf{z}_{x}^{j},\mathbf{z}_{s}^{k})].$ (16)
where $\tilde{y}=1$ denotes true samples and $\tilde{y}=0$ denotes fake samples.
@@
The optimizations of $f_{e}$, $f_{a}$ and $f_{x}$ using $\mathcal{L}_{total}$ and $g_{\omega}$ using $\mathcal{L}_{adv}$ are conducted adversarially. A well-trained disentangled graph encoder $f_{e}$, which effectively separates sensitive-irrelevant representations from node attributes and graph topology, can be adeptly employed in downstream tasks for enhancing fairness.
3.2 Graph Anomaly Detection